1.completing the square to solve 3x^2+10x=3.the students first step was to add 25 to each side of the equation,yielding 3x^2+10x+25=3+25.what should the student haven done first?



2.Eddie and Marco are working on a landing pad for their ultra light aircraft that will be 10 feet longer than a width. if the area of the landing pad is to be 1,200 square feet, find the length and width

Answer by stanbon(75887)

1.completing the square to solve 3x^2+10x=3.the students first step was to add 25 to each side of the equation,yielding 3x^2+10x+25=3+25.what should the student haven done first?
----
3x^2+10x = 3
Divide by 3 to get:
x^2 + (10/3)x = 1
Now, complete the square:
x^2 + (10/3)x + (5/3)^2 = 1 + (5/3)^2
(x+(5/3))^2 = 34/9
x + (5/2) = +-sqrt(34)/3
etc.
====================
2.Eddie and Marco are working on a landing pad for their ultra light aircraft that will be 10 feet longer than a width. if the area of the landing pad is to be 1,200 square feet, find the length and width.
----
Let width be "x" ft
Then length = "x+10" ft
---
Area = length*width
x(x+10) = 1200
x^2+10x-1200 = 0
Factor:
(x+40)(x-30) = 0
Positive solution:
x = 30 ft (width)
x+10 = 40 ft (length)
========================
Cheers,
Stan H.
==============================